OpenStudy (jamaljunior):

Which ordered pair is a solution to this equation? (x+7)y=10

OpenStudy (mathmale):

This is not a "linear function," because you have the product xy on the left side. Are you positive that you have copied down this problem correctly? Where are the answer choices?

OpenStudy (jamaljunior):

yes. https://gyazo.com/c25fbe758c50a3c7258f0c4b663424e7

OpenStudy (jamaljunior):

(x + 7) y = 10

OpenStudy (jungleman03):

Now all you have to do is substitute those in until you find your answer

OpenStudy (jungleman03):

for example (4,1) 4 would be x and 1 would be y

OpenStudy (jungleman03):

I also think your problem should be wriiten as this : (x+7) y=10 give yourself some room or it is considered incorrect like mathmale stated

OpenStudy (jamaljunior):

how do i substitute

OpenStudy (mathmale):

So, again, where are your four answer choices?

OpenStudy (jamaljunior):

https://gyazo.com/c25fbe758c50a3c7258f0c4b663424e7

OpenStudy (jungleman03):

(4,1) (3,1) (2,1) (13,2)

OpenStudy (mathmale):

Let's check out (13,2) (determine whether this is or is not a solution). In the given equation, let y=2 and x=13. Is the resulting equation true or false? Please show your work if you want feedback on it.

OpenStudy (skullpatrol):

Replace the letter with the corresponding number. $$\huge (x,y)$$

OpenStudy (jungleman03):

Skullpatrol summed it up :)

OpenStudy (mathmale):

Let's check out (13,2) (determine whether this is or is not a solution). In the given equation, let y=2 and x=13. Is the resulting equation true or false? Please show your work if you want feedback on it.

OpenStudy (jamaljunior):

how do i know if its true or not

OpenStudy (jamaljunior):

where do i start

OpenStudy (skullpatrol):

Are the two sides equal?

OpenStudy (mathmale):

would you please substitute as indicated: In the given equation, let y=2 and x=13. Is the resulting equation true or false? Please show your work if you want feedback on it. The resulting equation will either be true or it will not. 5 does not equal 4. 9 does equal 9. True

OpenStudy (mathmale):

Here's the equation: (x + 7) y = 10

OpenStudy (jamaljunior):

false

OpenStudy (skullpatrol):

Try the next pair.

OpenStudy (mathmale):

Please show your work, if you want feedback.

OpenStudy (jamaljunior):

well 7 does not = 10 so its false right?

OpenStudy (mathmale):

(x + 7) y = 10 becomes (13+7) ( 2 ) = 10 which in turn becomes (20)(2) = ? 10 This is what I'm asking you to do. Try the next pair, please.

OpenStudy (jamaljunior):

im so confused how does it before ( 2 ) = 10 and (20)(2)

OpenStudy (mathmale):

We have (13+7) ( 2 ) = 10 , which comes straight from (x+7)y = 10. I let x=13 and y=2. Then (13+7) = 20. Next, I multiply 20 by 2. Result? 40. 40 is not equal to 10, so (13,2) is not a solution of your (x+7)y=10.

OpenStudy (mathmale):

What is the next point? You need to determine whether the coordinates of this next point make the equation true or not.

OpenStudy (jamaljunior):

ok so (2+7) ( 1 ) = 10 9 x 1 = 9 so not true

OpenStudy (jamaljunior):

is that right

OpenStudy (jamaljunior):

i did it for (2,1)

OpenStudy (mathmale):

And your conclusion is correct: Not true, so (2,1) is NOT a solution. This leaves you with one more possible solution. What is it? Show your work as you check it in the equation.

OpenStudy (jamaljunior):

(3 + 7) (1) = 10 3 + 7 = 10 x 1 = 10

OpenStudy (jamaljunior):

so that one is true

OpenStudy (skullpatrol):

That fact makes it a "solution" to the equation.

OpenStudy (jamaljunior):

so thats right 3,1)

OpenStudy (mathmale):

Where did the '1' come from? Thought you had (x+7)y = 10. You were checking (3,1)? Important: Please do NOT use 'x' to denote multiplication. You could use " * " instead. x is read as a variable in algebra.

OpenStudy (mathmale):

So, what is your final answer to this question?

OpenStudy (jamaljunior):

(3, 1)